# Characterization of a fiber laser hydrophone for acoustic neutrino detection

### v1

@@ 1 Introduction
In the search for ultra-high energy neutrinos alternative detection techniques are being pursued. One of them is the detection of neutrino interactions through their acoustic signature in the deep sea [1, 2, 3]. This method benefits from the relatively long attenuation length of the sound in water, allowing to establishment of an acoustic neutrino telescope with a large detection volume with a relative sparse network of hydrophones. Still, for an acoustic telescope to observe the small predicted flux of ultra high energy neutrinos, the hydrophones are required to be sensitive enough to detect the weak signals induced by neutrino interactions. In addition, a neutrino telescope will consist of a large number of hydrophones.
Given the large number of hydrophones required for an acoustics neutrino telescope in combination with the required sensitivity, a promising technology is the fiber laser hydrophone technology as it combines sensitivity with a relative simple and cost-effective method of deployment. Fiber optic hydrophones (and fiber optic sensing) have been discussed in detail in literature. For scientific objectives additional requirements have to be met on the performance of the hydrophones, such as the operation in the deep-sea environment ($>1$ km deep), sensitivity (comparable to the lowest sea state noise levels, SS0-SS1) and large operating frequency bandwidth ($>10$ kHz). The anticipated acoustic signals vary with neutrino energy but are typically just above the sea state noise, with signal amplitudes ranging from 10 to several hundred millipascals [4, 5].
@@ 2 Hydrophone concept design
The concept of fiber hydrophone relies on the measurement of a wavelength shift in an optical fiber that is induced by sound pressure in water. A pressure wave will be converted to a wavelength shift by a transducer that is glued onto the fiber. The sensitivity of the hydrophone technology is determined by the performance of the transducer to efficiently convert pressure into strain in the optical fiber, but also of the optical fiber to convert strain into a wavelength shift. The wavelength shift is determined using a dedicated interferometer. The interferometer is a central part of the complete readout system. When designing and characterization of the hydrophone technology, as discussed in this paper, all three components have been investigated.
2.0.1 Transducer design
The transducer is in-house designed and is shown in figure 1. The design relies on one (or two) membranes placed on a cylindrical shaped housing. The hydrophone as shown in figure 1 is a cylindrical shaped transducer with a 250 $\mu$m thick membrane on the end of this cylinder. The dimensions of the membrane are determined by COMSOL and determines the position of the first mechanical eigenfrequency. The value of the eigenfrequency determines to first order the dynamical frequency range of the hydrophone. In the design a trade-off has to be made between the sensitivity and the dynamical range of the hydrophone.
A fiber is threaded through the membranes and glued. Particularly for this transducer, a hydrostatic pressure mechanism was introduced by an orifice in the housing, because the hydrophone is required to operate in the deep-sea and therefore to withstand high ambient pressure. The orifice allows water or oil to flow into the transducer to counterbalance the high pressure environment. Note however that a hydrophone with hydrostatic pressure compensation is expected to be less sensitive than an air-backed hydrophone. However, for an application in the deep-sea an air-backed volumes does not provide a solution, because the air is (completely) compressed by the ambient pressure.
@@ 2.0.1 Transducer design
The length of the transducer, i.e. 22 mm, is chosen to match the length of the fiber laser cavity (next section). The transducer is made of stainless steel, which is chosen because of its stiffness, density and ease of manufacturing. Then, two versions of the transducers have been produced: i) one membrane on one side of the cylinder as depicted in figure 1 and ii) with a membrane on each side on the cylinder. Fabrication and integration of the transducer was done at TNO. The membrane was laser-beam welded (LBW) on to the cylinder, while the orifices were EDM-drilled. The fiber was integrated in the transducer and glued on the membrane with glue (Loctite® EA9313). The fiber was not stripped from its coating before gluing. The pre-tension that was applied on the fiber amounted to about 0.1 N. Finally, to increase the robustness two end-caps are glued on either side of the cylinder to guide the fiber and to reduce the chance of breaking it.
2.1 Optical fiber hydrophone concept and signal formation
Dedicated optical fibers are realized that include a fiber laser as depicted in figure 2. The gain medium of the fiber laser is provided by the Erbium that is inserted, while the fiber Bragg grating (FBG) confines the laser cavity.
@@ 2.1 Optical fiber hydrophone concept and signal formation
For the discussion on strain sensitivity it is illustrative to discuss the strain dependence of the Bragg wavelength. As mentioned above the strain ($\varepsilon$) in the fiber induces a wavelength shift ($\Delta\lambda_{B}$) which can be described by [8]:
@@
where $p_{e}$ is an effective photo-elastic coefficient:
@@
Here $P_{ij}$ are the Pockels coefficients of the optical fiber, $n$ is the index of refraction and $\mu$ is the Poisson’s ratio. Given the values of $p_{e}\sim 0.22$ and $\lambda_{B}\sim 1550\;{\rm nm}$, the wavelength shift dependence on the strain is given by:
$\frac{\Delta\lambda_{B}}{\Delta\varepsilon}=1.2\left(\frac{pm}{\mu\varepsilon}\right)$ (3)
@@
$\phi=\frac{2\pi}{\lambda}\textsc{OPD}$ (4)
where $\lambda$ is the laser wavelength. A change in the central wavelength of the fiber laser induces a phase difference, which depends on the central wavelength and is given (for small wavelength changes) by:
$\frac{\Delta\phi}{\Delta\lambda}=\frac{2\pi}{\lambda^{2}}\textsc{OPD}$ (5)
The expressions as given in equations 5 and 3 are used to convert the measured phase to the strain in the fiber laser. The value OPD should be corrected for the index of refraction, i.e. the OPD is this case is $n$ times larger than the physical length of the difference in the armlength, where $n$ is the index of refraction of the glass fiber. In principle a larger OPD provides a higher sensitivity, but large armlengths in the interferometer are unpracticle and susceptible for additional (internal fiber) noise. It was found that in our experimental set-up the optimal optical path difference was between 4 and 10 meter.
The fiber laser consists of a 22 mm long Erbium-doped Distributed Feedback Laser (DFB) produced by Exail. The length of the fiber laser is important and should be as short as possible to induce the maximum strain for the same pressure wave. Furthermore, the length of the transducer housing is matched to the length of the fiber laser.
@@ 3.1 Interrogator
Secondly, an interferometer is used to determine the wavelength shift of the fiber laser light. We use a custom built interferometer for this based on a standard 3x3 coupler. While the input of the interferometer consists of three inputs, we only used two arms in the interferometer. A fixed and predetermined optical path difference (OPD) in the two arms has been chosen. The output of the 3x3 coupler consists of three light intensities ($I_{x}$) as indicated in figure 3. The outputs of the three arms are identical, but each is shifted in phase over $\frac{2}{3}\pi$. In addition, each intensity depends on the optical path difference as defined in the equation 3.1.
@@
Because the optical path difference (OPD) is known, the wavelength shift of each hydrophone in the string can be accurately determined when the three output intensities are measured.
Then, thirdly, we convert the 3 optical signals from the interferometer arms into a digital signal. To do so we collect the light intensities from the interferometer using standard commercial available photo-diodes and converted using a 24-bit ADC at a sample rate of 200 kHz. Finally, we use the same DAQ system to register the signals from a commercial available reference hydrophone (section 3.3).
@@ 3.3 Acoustic setup
To characterize the fiber hydrophone, the experimental setup and the hydrophones as described above were taken to the anechoic basin at the TNO laboratories. The experimental set-up in the basin is depicted schematically in figure 5. The basin it self measures 8x8x10 m3.
@@
A mechanical support structure was produced to which both the fiber hydrophone and the reference hydrophone are attached.
@@ 3.4 Pressure calibration
The calibration of the fiber optic hydrophone is established using the reference hydrophone, i.e a commercial available Bruel & Kjaer BK8103. In turn this hydrophone is calibrated using a pistonphone (type BK4229 with coupler UA 0548), which provides a well-defined signal at 250 Hz to the coupled reference hydrophone. This signal has been calibrated and provides a well defined signal of $166\;{\rm dB\;re}\;\mu{\rm Pa}$ to the reference hydrophone. The response of the BK8103 is specified to be flat up to 50 kHz, so that the calibration of the reference hydrophone is valid for the range at which the fiber optic hydrophone is tested.
@@ 4.1 Hydrophones under test
To assess the performance of the hydrophones a number of different configurations have been integrated and characterized. In table 1 the various hydrophone configurations are listed. Each of the hydrophones have been characterized, but in addition we performed a number of measurements to better understand the impact of several issues during integration. One hydrophone has been subjected to a qualification measurement.
@@
All hydrophones were filled with water. To properly fill the volumes within the transducers, an experimental set up was made to fill the transducers without air bubbels. After integration the transducers were put in a reservoir under vacuum and water was led in the reservoir. This controlled way of filling ensured that the amount of air in the volumes after being filled was minimal.
@@ 4.2 Noise
Prior to the integration of the fiber laser hydrophones, bare fiber lasers coupled to the interrogatore have been characterized. Using the interrogator the phase noise of the fiber laser without an transducer attached to it has been determined in an acoustically insulated environment and free from vibrations. In figure 6 the wavelength noise for one of the fiber lasers is shown for as a function of the frequency. In comparison we show the noise measurement of an acoustic insulated fiber laser. While the packaged hydrophone was read out using an interferometer with a OPD of 4 m, the measurement on the fiber laser was carried with an interferometer with an OPD of 10 m. In addition, the fiber laser was kept in an acoustically isolated block of aluminium, free from vibrations. Hence, the measurements on the bare fiber laser presents the noise floor of the interrogator. As can be seen from the figure, the noise level of the integrated and the bare fibers do not differ much at higher frequencies, i.e. above 1 kHz. At a frequency of 10 kHz, the difference in the noise level is about 2 dB. At lower frequencies the integrated fiber pick up acoutic and vibration noise. Spikes that appear in the noise spectra orginate from electromagnetic interference.
Figure 6: Power spectral density of the phase noise of a fiber laser integrated in an optical fiber hydrophone compared to the noise of an acoustically insulated bare fiber laser.
@@ 4.3 Time trace and Bode diagram
Figure 7: Example spectrogram of a single membrane fiber hydrophone (top panel) and time traces from the fiber and reference hydrophones (bottom panel) in a frequency sweep measurement.
In the lower panel of figure 7 the time trace for both the fiber hydrophone and the reference hydrophones are shown. Here, we present the uncalibrated data to demonstrate that (i) the output of the projector, as recorded by the reference, is not flat across the spectrum, and (ii) the transfer function of the fiber hydrophone is also not uniform, but rather complex. It shows the importance of the reference hydrophones in these measurements to correct for the distortions in the source and the investigae the transfer function of the fiber hydrophone.
Frequency sweeps were used to determine transfer functions and were plotted in general Bode diagrams consisting of a Bode magnitude plot and a Bode phase plot. In addition we determine the coherence between the signals in the fiber hydrophone and the reference hydrophone. The coherence confirms whether the measured signals is observed in both hydrophones and hence wether the input is of acoustic origin. We compute (magnitude of the) transfer function ($T_{xy}$) and the coherence ($C_{xy}$) according:
@@
Figure 8: Response of a single membrane hydrophone to a sweep signal with a frequency between 200 Hz and 20 kHz. In the top panel the Bode magnitude plot is shown; in the middle panel the Bode phase plot is shown, while the bottom panel shows the coherence of the signal between the fiber optic hydrophone and the reference hydrophone.
In figure 8 we show the Bode plot of a frequency response measurements in which an acoustic frequency sweep between 200 Hz and 20 kHz was offered to both the (double membrane) fiber hydrophone and the reference hydrophone. The panels in this figure show (from top to bottom) the magnitude and the phase of the transfer function as well as the coherence (cross spectral density) between the fiber hydrophone and the reference. The coherence in the frequency range of 200 Hz - 20 kHz is on average high, albeit that there are several dips, which can be attributed to non-acoustic signals such as electronic noises.
@@
Bode plots are used to determine the strain sensitivity of the transducer. From the measured laser phase shift (and hence wavelength shift) the strain sensitivity can be determined using equation 3. In figure 9 we show the strain sensitivity for both the single and double membrane hydrophones. The sensitivity is given for the frequency range in which a frequency sweep has been carried out and in which the coherence is sufficient (i.e. $C_{xy}>0.7$). As we characterized two single membrane transducers, we compared their transfer functions, and no significant differences were found.
In addition the sensitivity as predicted by the COMSOL finite element analysis is given. As can be seen from the figure, the measured strain sensitivity matches reasonably with what has been determined using the design software, albeit that the position of the resonance peak is found at lower frequencies than what has been designed.
Figure 9: Measured and designed strain sensitivity of a single and double membrane hydrophone as a function of the frequency.

### v2

@@ 1 Introduction
In the search for ultra-high-energy neutrinos alternative detection techniques are being pursued. One of them is the detection of neutrino interactions through their acoustic signature in the deep sea [1, 2, 3]. This method benefits from the relatively long attenuation length of the sound in water, allowing the establishment of an acoustic neutrino telescope with a large detection volume with a relatively sparse network of hydrophones. Still, for an acoustic telescope to observe the small predicted flux of ultra-high-energy neutrinos, the hydrophones are required to be sensitive enough to detect the weak signals induced by neutrino interactions. Additionally, achieving an effective detection volume of approximately 100 km3 or larger would necessitate a large-scale array, comprising several thousand hydrophones [4].
Given the large number of hydrophones required for an acoustic neutrino telescope in combination with the required sensitivity, a promising technology is the fiber laser hydrophone technology as it combines sensitivity with a relatively simple and cost-effective method of deployment. Fiber optic hydrophones have been discussed in detail in literature, see e.g. [5, 6]. For the application in an acoustic neutrino telescope, additional requirements have to be met on the performance of the hydrophones, such as the operation in the deep-sea environment ($>1$ km deep), sensitivity (comparable to the lowest sea state noise levels, SS0-SS1) and a large operating frequency bandwidth ($>10$ kHz). The expected acoustic signals depend on neutrino energy but generally remain just above the sea state noise. For neutrino energies between $10^{10}$ and $10^{12}$ GeV, the signal amplitudes range from approximately 10 to several hundred millipascals at a distance of 1 km [7, 8].
@@ 2 Hydrophone concept design
The concept of fiber hydrophones relies on the measurement of a wavelength shift in an optical fiber that is induced by sound pressure in water. A pressure wave will be converted to a wavelength shift by a transducer that is attached to the fiber. The sensitivity of this hydrophone technology depends on both the transducer’s ability to efficiently convert pressure into strain in the optical fiber and the fiber’s ability to translate that strain into a wavelength shift. This shift is then measured using a custom designed interferometer, which is a key component of the readout system. In this paper, we thoroughly examine and discuss all three essential elements: the transducer, the optical fiber, and the read-out system (or interrogator).
2.1 Transducer design
The transducer is in-house designed and is shown in figure 1. The design relies on one (or two) membranes placed on a cylindrical shaped housing. The optical fiber is threaded through the center of the membranes and subsequently glued. The dimensions of the membrane, i.e., a thickness of 250 $\mu$m and a radius of 5 mm, were determined using COMSOL and define the mechanical response of the transducer. The eigenfrequency, which corresponds to the position of the mechanical resonance peak, is particularly important, as the transducer’s sensitivity is enhanced at this peak and declines sharply at higher frequencies.
@@ 2.1 Transducer design
The length of the transducer, i.e. 22 mm, is chosen to match the length of the fiber laser cavity (next section). The transducer is made of stainless steel, which is chosen because of its stiffness, density and ease of manufacturing. Then, two versions of the transducers have been produced: i) one membrane on one side of the cylinder as depicted in figure 1 and ii) with a membrane on each side on the cylinder. Fabrication and integration of the transducer was done at TNO. The membrane was laser-beam welded (LBW) on to the cylinder, while the orifices were EDM-drilled. The fiber was integrated in the transducer and glued on the membrane with glue (Loctite® EA9313). The fiber was not stripped from its coating before gluing. The pre-tension that was applied on the fiber amounted to about 0.1 N. Notably, the strength of the glue remained unchanged over time. After more than three years in storage, we remeasured the hydrophone transducer’s response and found no detectable difference. Finally, to increase the robustness two end-caps are glued on either side of the cylinder to guide the fiber and to reduce the chance of breaking it.
Particularly for this transducer, a hydrostatic pressure mechanism was introduced by an orifice in the housing, because the hydrophone is required to operate in the deep-sea and therefore to withstand high ambient pressure. The orifice allows water or oil to flow into the transducer to counterbalance the high pressure environment. Note however that a hydrophone with hydrostatic pressure compensation is expected to be less sensitive than an air-backed hydrophone. However, for an application in the deep-sea an air-backed volumes does not provide a solution, because the air is (completely) compressed by the ambient pressure.
2.2 Fiber laser hydrophone concept and signal formation
Dedicated optical fibers are realized that include a fiber laser as depicted in figure 2 [11, 12]. The gain medium of the fiber laser is provided by the Erbium-doped fiber core, while the fiber Bragg grating (FBG) confines the laser cavity.
@@ 2.2 Fiber laser hydrophone concept and signal formation
For the discussion on the strain gauge factor it is illustrative to discuss the strain dependence of the Bragg wavelength. As mentioned above the strain ($\varepsilon$) in the fiber induces a wavelength shift ($\Delta\lambda_{B}$) which can be described by [13]:
@@
where $\lambda_{B}$ is the central laser wavelength and $p_{e}$ is an effective photo-elastic coefficient:
@@
Here $P_{ij}$ are the Pockels coefficients of the optical fiber, $n$ is the index of refraction and $\mu$ is the Poisson’s ratio [14]. Given the values of $p_{e}\sim 0.22$ and $\lambda_{B}\sim 1550\;{\rm nm}$, the wavelength shift dependence on the strain is given by:
$\frac{\Delta\lambda_{B}}{\varepsilon}=1.2\left(\frac{pm}{\mu\varepsilon}\right)$ (3)
@@
$\phi=\frac{2\pi}{\lambda_{B}}\textsc{OPD}$ (4)
A change in the central wavelength of the fiber laser induces a phase difference, which depends on the central wavelength and is given (for small wavelength changes) by:
$\frac{\Delta\phi}{\Delta\lambda_{B}}=\frac{2\pi}{\lambda_{B}^{2}}\textsc{OPD}$ (5)
The expressions as given in equations 5 and 3 are used to convert the measured phase to the strain in the fiber laser. The value OPD should be corrected for the index of refraction, i.e. the OPD is this case is $n$ times larger than the physical length of the difference in the arm length, where $n$ is the index of refraction of the glass fiber. In principle a larger OPD provides a higher sensitivity, but large arm lengths in the interferometer are unpractical and susceptible for additional (internal fiber) noise. It was found that in our experimental set-up the optimal optical path difference was between 4 and 10 meter.
@@ 3.1 Interrogator
Secondly, an interferometer is used to determine the wavelength shift of the fiber laser light. We use a custom built interferometer for this based on a standard 3x3 fiber coupler. While the input of the interferometer consists of one input and three outputs, we only used two arms in the interferometer. A fixed and predetermined optical path difference (OPD) in the two arms has been chosen. The output of the 3x3 coupler consists of three light intensities ($I_{x}$) as indicated in figure 3. The outputs of the three arms are identical, but each is shifted in phase over $\frac{2}{3}\pi$. In addition, each intensity depends on the optical path difference as defined in the equation 3.1.
@@
Because the optical path difference (OPD) is known, the wavelength shift can be accurately determined when the three output intensities are measured.
Then, thirdly, we convert the three optical signals from the interferometer arms into a digital signal. To achieve this, we collect the light intensities from the interferometer using commercially available photo-diodes, which are then converted using a 24-bit ADC at a sample rate of 200 kHz. Finally, we use the same DAQ system to register the signals from a commercially available reference hydrophone (section 3.3).
@@ 3.3 Acoustic setup
To characterize the fiber hydrophone, the experimental setup as described above was taken to the anechoic basin at the TNO laboratories. The experimental set-up in the basin is depicted schematically in figure 5. The basin itself measures 8x8x10 m3.
@@
A mechanical support structure was produced to which both the fiber hydrophone and the reference hydrophone were attached.
@@ 3.4 Pressure calibration
The calibration of the fiber hydrophone is established using the reference hydrophone, i.e a commercial available Bruel & Kjaer BK8103. In turn this hydrophone is calibrated using a pistonphone (type BK4229 with coupler UA 0548), which provides a well-defined signal at 250 Hz to the coupled reference hydrophone. This signal has been calibrated and provides a well defined signal of $166\;{\rm dB\;re}\;\mu{\rm Pa}$ to the reference hydrophone. The response of the BK8103 is specified to be flat up to 50 kHz, so that the calibration of the reference hydrophone is valid for the range at which the fiber hydrophone is tested. Both the reference hydrophone and the pistonphone undergo periodic control measurements. Moreover, cross measurements with various reference hydrophones have been carried out for additional verification of the hydrophone’s response.
@@ 4.1 Hydrophones under test
To assess the performance of the fiber hydrophones a number of different configurations have been integrated and characterized. In table 1 the various fiber hydrophone configurations are listed. Each of the fiber hydrophones have been characterized, but in addition we performed a number of measurements to better understand the impact of several issues during integration. One fiber hydrophone has been subjected to a pressure qualification measurement.
@@
All fiber hydrophones were filled with water. To properly fill the volumes within the transducers, an experimental set up was made to fill the transducers without air bubbles. After integration the transducers were put in a reservoir under vacuum and water was led in the reservoir. This controlled way of filling ensured that the amount of air in the volumes after being filled was minimal.
@@ 4.2 Noise
Prior to integrating the fiber hydrophones, bare fiber lasers coupled to the interrogator were characterized. Using the interrogator the phase noise of the fiber laser without an attached transducer was measured. The fiber laser was kept in an acoustically isolated block of aluminium. In addition the interrogator was kept free from vibrations. The set-up used included an interferometer with a optical path length (OPD) of 10 m. Hence, the measurements on the bare fiber laser presents the noise floor of the interrogator. In figure 6 the wavelength noise - obtained from the phase shift using equation 5 - for one of the bare fiber lasers is shown for as a function of the frequency. For comparison we show the noise measurement on a packaged hydrophone both in water and in air. Here a set-up that included an interferometer with an OPD of 4 m as used. As can be seen from the figure, the noise level of the integrated and the bare fibers do not differ much at higher frequencies, i.e. above 1 kHz. At a frequency of 10 kHz, the difference in the noise level is about 2 dB. At lower frequencies, the integrated fiber picks up both acoustic and vibration noise. Peaks in the noise spectrum (clearly seen in the fiber laser spectrum) originate from electromagnetic interference, as some noise is coupled into the analog front-end of the interrogator.
Figure 6: Power spectral density of the phase noise of a fiber laser integrated in a fiber hydrophone compared to the noise of an acoustically insulated bare fiber laser.
@@ 4.3 Time trace and Bode diagram
Figure 7: Example spectrogram of a single membrane fiber hydrophone (top panel). The colors in this plot qualitatively represent the increasing signal amplitude, transitioning from green to purple. In the bottom panel the time traces from the fiber and reference hydrophones in a frequency sweep measurement are shown.
In the lower panel of figure 7 the time trace for both the fiber hydrophone and the reference hydrophones are shown. Here, we present the uncalibrated data to demonstrate that (i) the output of the projector, as recorded by the reference, is not flat across the spectrum, and (ii) the transfer function of the fiber hydrophone is also not uniform, but rather complex. It shows the importance of the reference hydrophones in these measurements to correct for the distortions in the source and the investigate the transfer function of the fiber hydrophone.
Frequency sweeps were used to determine transfer functions and were plotted in general Bode diagrams consisting of a Bode magnitude plot and a Bode phase plot. In addition we determine the coherence between the signals in the fiber hydrophone and the reference hydrophone. The coherence confirms whether the measured signal is observed in both hydrophones and hence whether the input is of acoustic origin. We compute (magnitude of the) transfer function ($T_{xy}$) and the coherence ($C_{xy}$) according:
@@
Figure 8: Response of a single membrane hydrophone to a sweep signal with a frequency between 200 Hz and 20 kHz. In the top panel the Bode magnitude plot is shown; in the middle panel the Bode phase plot is shown, while the bottom panel shows the coherence of the signal between the fiber hydrophone and the reference hydrophone.
In figure 8 we show the Bode plot of a frequency response measurements in which an acoustic frequency sweep between 200 Hz and 20 kHz was offered to both the (single membrane) fiber hydrophone and the reference hydrophone. The panels in this figure show (from top to bottom) the magnitude and the phase of the transfer function as well as the coherence (cross spectral density) between the fiber hydrophone and the reference. The coherence in the frequency range of 200 Hz - 20 kHz is on average high, albeit that there are several dips, which can be attributed to non-acoustic signals such as electronic noises.
@@
Bode plots are used to determine the strain gauge factor of the transducer. From the measured laser phase shift (and hence wavelength shift) the strain gauge factor can be determined using equation 3. In figure 9 we show the strain gauge factor for both the single and double membrane fiber hydrophones. The gauge factor is given for the frequency range in which a frequency sweep has been carried out and in which the coherence is sufficient (i.e. $C_{xy}>0.7$). As we characterized two single membrane transducers, we compared their transfer functions, and no significant differences were found.
In addition, the strain gauge factor as predicted by the COMSOL finite element analysis is given. As can be seen from the figure, the measured strain gauge factor reasonably matches the valued determined using the design software, albeit that the position of the resonance peak is found at lower frequencies than originally designed.
Figure 9: Measured and designed strain gauge factor of a single and double membrane fiber hydrophone as a function of the frequency.